Circular No. 7178

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 K5 (LINEAR) Another apparently asteroidal object discovered by LINEAR (discov-

ery observation below; additional astrometry on MPEC 1999-K33), and posted on The NEO Confirmation Page, has been reported as cometary by several observers.

1999 UT  α_{2000}  δ_{2000}  m_{2}
May 20.31735  19^{h}44^{m}23.^{s}12  −9^{o}15′02′′.8  17.2

CCD frames taken by D. D. Balam (Victoria) on May 23 show a condensed coma with a 16′′ fan-shaped tail in p.a. 303^{o}. On May 24, L. Kornoš and P. Kolény (Modra) report a coma diameter of ∼ 15′′ and a short tail in p.a. 330^{o}, and G. Hug (Farpoint Observatory) indicated a hint of coma in p.a. ∼ 300^{o}. Preliminary parabolic orbital elements by B. G. Marsden:

T = 2000 July 4.806 TT  ω = 241.^{o}142
Ω = 106.385 }2000.0
q = 3.27508 AU  i = 89.495
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 22 19^{h}43.^{m}74 − 9^{o}24.^{′}6 4.331 4.963 123.^{o}5 9.^{o}8 16.1
June 1 19 38.98 −10 28.1 4.137 4.899 134.3 8.5 16.0
11 19 32.61 −11 43.2 3.966 4.836 145.4 6.8 15.8
21 19 24.74 −13 09.6 3.823 4.773 156.7 4.8 15.7
July 1 19 15.61 −14 46.0 3.713 4.711 167.8 2.6 15.6
11 19 05.58 −16 29.7 3.638 4.650 173.5 1.4 15.5
21 18 55.16 −18 17.4 3.601 4.589 164.6 3.4 15.4

SUPERNOVA 1999cf IN UGC 8539 M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag ∼ 15.9) on an unfiltered CCD frame (limiting mag ∼ 17) taken with the 0.35-m Tenagra I automated supernova patrol telescope at Sonoita, AZ, on May 22.27 UT, with a confirming observation on May 24.13. SN 1999cf is located at α = 13^{h}33^{m}29^{s}.71, δ = +33^{o}03′22′′.9 (equinox 2000.0), which is ∼ 7′′.9 east and 49′′.7 north of the center of UGC 8539. The new object does not appear on either the first Palomar Sky Survey digital image or on a CCD frame taken with the same equipment on 1998 June 28.28 (limiting mag 18.5).

1999 May 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green